OpenStudy (anonymous):

A box contains 12 transistors, 3 of which are defective. If 3 are selected at random, find the probability? All are Defective = None are defective= I came up with answers I dont think i am correct thank you All = 1/288 None = 7/24

OpenStudy (anonymous):

A box contains 30 transistors, 6 of which are defective. If 6 are selected at random find the probability that a) all are defective. The probability that the 1st selection will be defective is 6/30, the probability that the 2nd selection will be defective is 5/29, the probability that the 3rd selection will be defective is 4/28, etc., the probability that the 6th selection will be defective is 1/25. Multiply those probabilities: 6/30 * 5/29 * 4/28 * 3/27 * 2/26 * 1/25 = 720 / 427518000 = 1 / 593775.

OpenStudy (anonymous):

similar problem ^

OpenStudy (anonymous):

But how do you find probability none defective? i got this one right, but i am struggling with none defective